# DataChef: Cooking Up Optimal Data Recipes for LLM Adaptation via Reinforcement Learning

### v1

@@ Abstract
In the current landscape of Large Language Models (LLMs), the curation of large-scale, high-quality training data is a primary driver of model performance. A key lever is the data recipe, which comprises a data processing pipeline to transform raw sources into training corpora. Despite the growing use of LLMs to automate individual data processing steps, such as data synthesis and filtering, the overall design of data recipes remains largely manual and labor-intensive, requiring substantial human expertise and iteration. To bridge this gap, we formulate end-to-end data recipe generation for LLM adaptation. Given a target benchmark and a pool of available data sources, a model is required to output a complete data recipe that adapts a base LLM to the target task. We present DataChef-32B, which performs online reinforcement learning using a proxy reward that predicts downstream performance for candidate recipes. Across six held-out tasks, DataChef-32B produces practical recipes that reach comparable downstream performance to those curated by human experts. Notably, the recipe from DataChef-32B adapts Qwen3-1.7B-Base to the math domain, achieving 66.7 on AIME’25 and surpassing Qwen3-1.7B. This work sheds new light on automating LLM training and developing self-evolving AI systems.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) DeepSeek-AI et al. (2025); OpenAI (2025) has underscored the central role of data in determining model capabilities, making a shift toward data-centric AI Jakubik et al. (2024). The composition and quality of training data emerge as decisive factors in shaping downstream performance. Practically, constructing effective training data requires a well-designed multi-stage pipeline that processes heterogeneous raw data through a sequence of operations, such as transformation, filtering, mixing, synthesis, and refinement, tailored to specific training goals or stages Yang et al. (2025); Cai et al. (2025). In this work, we formalize the concept of a data recipe, defined as the combination of the processing pipeline and the resulting training data.
@@
In practice, constructing a data recipe typically involves substantial human expertise and manual effort, with human experts orchestrating a subset of data processing operations in a specific order Penedo et al. (2025); Gururajan et al. (2024). While LLMs are widely used in individual processing operations, such as data filtering Liu et al. (2024), data selection Zhang et al. (2025d), and data synthesis Mitra et al. (2024); Huang et al. (2024), they still follow a human-designed prompt or pattern to prepare data. Recent studies have explored automating data pipeline orchestration to reduce the reliance on manual effort. In particular, Data-Juicer Sandbox Chen et al. (2025a) proposes a Probe-Analyze-Refine workflow to identify the most impactful operators from a predefined operation pool, combine effective operations, and optimize data utilization through systematic experiments in data processing, model training, and evaluation with model performance as feedback. However, the continuous scaling of data and model sizes, coupled with the increasing complexity of processing operations, renders an exhaustive exploration of the combinatorial space of data recipes infeasible. Therefore, an essential question arises: can AI systems automatically generate a data recipe for training LLMs, including the orchestration of data pipelines and the implementation of each operation, in a cost-efficient way?
@@
To bridge this gap, we introduce a new task: end-to-end data recipe generation for LLM adaptation. As shown in Fig 1(a), given a target benchmark and a pool of available data sources, the objective is to generate a complete data recipe by specifying the precise data processing pipeline to yield training data for adapting an LLM to the target task. Producing an effective data recipe requires strong reasoning abilities, as it involves analyzing heterogeneous data sources, applying domain-specific processing operations, and generating corresponding executable code. While recent studies have demonstrated the effectiveness of reinforcement learning in enhancing LLM reasoning abilities in complex domains, such as coding Zeng et al. (2025) and mathematical reasoning Yeo et al. (2025), applying this paradigm to our task poses two key challenges: (1) Data absence: end-to-end data recipe generation is a previously unexplored task, for which no curated datasets or standardized evaluation benchmarks exist. (2) Expensive and delayed supervision: while downstream performance naturally serves as the reward signal, it is impractical to directly incorporate downstream LLM training into an online reinforcement learning loop.
@@
To address these challenges, we curate a comprehensive task pool comprising 31 widely used benchmarks across 10 distinct domains. These domains encompass reasoning-heavy fields, such as mathematics and coding, as well as knowledge-centric fields, such as finance, medicine, and the natural sciences. The pool is partitioned into 25 training tasks and 6 held-out evaluation tasks, with each task supported by 8–15 source training datasets. We further propose a data verifier that assesses training data quality directly without performing model training, providing a low-cost, instant reward signal in online RL. We systematically validate that the data verifier prediction correlates well with downstream model performance across domains. Leveraging cold-start fine-tuning and online reinforcement learning, we present DataChef-32B, an LLM specialized for generating optimal data recipes.
Through extensive evaluations, DataChef-32B demonstrates matching capability to Gemini-3-Pro on data recipe generation. Fig. 1(b) shows that DataChef-32B produces effective data recipes on 6 hold-out tasks, yielding high performance on downstream benchmarks. The generated recipes outperform the best individual data source on most tasks. Notably, on the math and atmosphere domain, recipes from DataChef-32B adapt Qwen3-1.7B-Base to achieve 66.7 on AIME’25 AIME (2025) and 46.3 on ClimaQA Manivannan et al. (2025), respectively, surpassing Qwen3-1.7B with industry-level post-training on expert-curated data recipes.
@@ 2. $\bullet$
    We construct a large-scale and diverse data pool covering 19 domains, 31 benchmarks, and 257 datasets to support this task.
@@ 3. $\bullet$
    We propose an efficient learning framework with a proxy reward that enables scalable online RL. Extensive experiments show that our DataChef-32B achieves performance comparable to that of top-tier proprietary models on the data recipe generation task.
@@ 3 Methodology
Figure 2: Illustration of DataChef training framework. Given a task, a policy LLM generates a data recipe, which is executed to produce a training dataset. The Data Verifier then evaluates a sampled subset to provide a scalar reward, guiding the policy update via GRPO to optimize for data quality and executability.
@@ 3.2 Task Pool Construction
Seed Task Curation. As a previously unexplored task, data recipe generation lacks a canonical corpus. To bridge this gap, we construct a diverse task pool encompassing 19 heterogeneous domains, including reasoning, coding, and knowledge-intensive fields such as healthcare, finance, and natural science. For each domain, we select representative benchmarks (e.g., GSM8K and AIME’25 for mathematics), totaling 31 benchmarks. For each benchmark, we retrieve relevant candidate datasets from Hugging Face, prioritizing those with high community engagement (downloads and likes), yielding a repository of 257 distinct data sources. From this collection, we construct 25 seed tasks for training and reserve 6 held-out tasks (3 in-domain, 3 out-of-domain) for evaluation. Comprehensive details of the selected benchmarks are provided in Appx. A.1.
@@ 3.3 End-to-end Data Recipe Generation
Cold-start Initialization. Training the policy from scratch using RL is non-trivial due to the low executability of data recipes, leading to sparse, high-variance rewards and ineffective exploration Shao et al. (2024); Liu et al. (2025c). To mitigate this, we employ a cold-start Supervised Fine-Tuning (SFT) phase. We observe that decoupling reasoning and coding yields superior inference-time performance (as discussed in Sec. 4.4). Therefore, we construct a high-quality demonstration set using a decoupled generation process: a strong reasoning model proposes plans, and a specialized coding model implements them. We filter these pairs by execution success and data quality, retaining only valid recipes. Initializing $\pi_{\phi}$ on this curated dataset equips the policy with a foundational capability for code generation, significantly stabilizing the subsequent RL phase.
Table 1: Main Results on six held-out tasks. We report the mean Data Verifier Score $\mathrm{DVS}_{\mathrm{avg}@32}$ and the Downstream Benchmark Score $\mathrm{DBS}$, where the Average column presents $\mathrm{DBS}$ as a normalized percentage relative to Sourcebest (100.0). Qwen3-Next $\oplus$ Kimi-K2 denotes a combination using Qwen3-Next-80B for reasoning and Kimi-K2-Instruct for coding. DataChef-32B achieves performance comparable to the closed-source Gemini-3-Pro and significantly outperforms other open-source baselines across all settings.
@@
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$
Sourceavg - 6.1 - 23.4 - 6.3 - 63.9
Sourcebest - 8.5 - 39.6 - 10.3 - 100.0
Expert - 20.8 - 33.3 - 25.7 - 123.3
@@
Qwen3-Next $\oplus$ Kimi-K2 48.7 8.9 78.3 23.3 39.2 7.4 55.4 78.6
@@
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$
Sourceavg - 26.0 - 41.7 - 49.3 - 65.1
Sourcebest - 43.6 - 63.7 - 70.3 - 100.0
Expert - 44.2 - 73.4 - 59.8 - 101.0
@@
Qwen3-Next $\oplus$ Kimi-K2 41.5 42.6 54.7 64.0 8.6 4.1 34.9 68.0
@@ 5. $\bullet$
where $\lambda_{\emptyset}$ and $\lambda_{\mathrm{fmt}}$ are positive penalty coefficients. Please refer to Appx. A.2 for a detailed description of the category definitions used in the prompt.
@@ 4.1 Setups
Figure 3: Correlation analysis of data evaluation metrics. (left) We summarize the Pearson correlation coefficients across all six evaluated tasks. (right) We detail the relationship between metric scores (X-axis) and downstream performance (Y-axis) on Language and Code tasks. The Data Verifier maintains a strong, consistent positive correlation across disparate domains. Please refer to Fig. 8 in Appx. D for complete results.
@@
Baselines. We compare DataChef-32B against three categories of models: (1) parameter-matched model: Qwen3-32B; (2) open-source flagships: Kimi-K2-Instruct Team et al. (2025) and Qwen3-Next-80B-A3B-Thinking; (3) closed-source SOTA: Gemini-3-Pro Google (2025). Additionally, we incorporate the following results as reference: First, to benchmark the raw data quality, we manually format each available source and report the average ($\textsc{Source}_{\mathrm{avg}}$) and best ($\textsc{Source}_{\mathrm{best}}$) downstream performance among these single-source datasets. Second, as a high-standard reference, we report Expert: the performance of Qwen3-1.7B optimized via industry-grade post-training with expert-curated recipes. Detailed experiment setups are provided in Appx. B.
@@ 4.2 Main Results
Main Comparison. Table 1 presents the performance of DataChef-32B against baselines across in-domain and out-of-domain tasks. DataChef-32B achieves superior performance compared to a strong practical baseline, Qwen3-Next $\oplus$ Kimi-K2, which leverages open-source state-of-the-art specialized models (Qwen3-Next-80B-A3B-Thinking for reasoning and Kimi-K2-Instruct for coding). Specifically, our end-to-end model surpasses this composite system with average improvements of +8.6% and +9.2% in $\mathrm{DVS}_{\mathrm{avg}@32}$, and +10.7% and +7.4% in $\mathrm{DBS}$ on in-domain and out-of-domain tasks, respectively. Notably, DataChef-32B achieves performance comparable to the closed-source top-tier Gemini-3-Pro, demonstrating exceptional robustness and effectiveness in automated data recipe generation.
Surpassing Human Baselines. By selecting the most promising recipe from 32 samples (Oracle Upper Bound), DataChef-32B outputforms $\textsc{Source}_{\mathrm{best}}$ on most tasks, achieving an average score of 130.3 on in-domain benchmarks. This indicates that DataChef goes beyond simple dataset selection and synthesizes novel data processing pipelines, including effective selection, mixing, synthesis, and filtering, which are superior to raw manual formatting. Remarkably, it achieves 66.7 on AIME’25 and 46.3 on ClimaQA, even surpassing the Expert baseline with industry-level post-training on expert-curated data recipes. These results underscore the potential of fully automating data recipe generation for LLM training.

### v2

@@ Abstract
In the current landscape of Large Language Models (LLMs), the curation of large-scale, high-quality training data is a primary driver of model performance. A key lever is the data recipe, which comprises a data processing pipeline to transform raw sources into training corpora. Despite the growing use of LLMs to automate individual data processing steps, such as data synthesis and filtering, the overall design of data recipes remains largely manual and labor-intensive, requiring substantial human expertise and iteration. To bridge this gap, we formulate end-to-end data recipe generation for LLM adaptation. Given a target benchmark and a pool of available data sources, a model is required to output a complete data recipe that adapts a base LLM to the target task. We present DataChef-32B, which performs online reinforcement learning using a proxy reward that predicts downstream performance for candidate recipes. Across six held-out tasks, DataChef-32B produces recipes that yield performance comparable to those curated by human experts. Notably, the recipe from DataChef-32B adapts Qwen3-1.7B-Base to the math domain, achieving 66.7 on AIME’25 and surpassing the official post-training checkpoint (Qwen3-1.7B). This work sheds new light on automating LLM training and developing self-evolving AI systems.
@@ 1 Introduction
The rapid evolution of LLMs DeepSeek-AI et al. (2025); OpenAI (2025) has precipitated a shift toward data-centric AI Jakubik et al. (2024), identifying the composition and quality of training data as decisive factors in shaping model performance. In practice, constructing effective training data requires a well-designed multi-stage pipeline that processes heterogeneous raw data through a sequence of operations, such as transformation, filtering, mixing, synthesis, and refinement, tailored to specific training goals or stages Yang et al. (2025); Cai et al. (2025). In this work, we formalize such processing pipelines with the resulting dataset as a data recipe.
@@
Currently, formulating an effective data recipe relies heavily on human heuristics, where experts manually orchestrate data processing operations and iteratively refine them based on empirical feedback Penedo et al. (2025); Gururajan et al. (2024). While LLMs are widely used to automate individual pipeline components, such as data filtering Liu et al. (2024); Zhang et al. (2025d) and synthesis Mitra et al. (2024); Huang et al. (2024), they typically operate under rigid, hand-crafted prompts or algorithms. Recent studies have explored automating data pipeline orchestration to reduce the reliance on manual effort. In particular, Data-Juicer Sandbox Chen et al. (2025a) proposes a Probe-Analyze-Refine workflow to identify the most impactful operators from a predefined operation pool, combine effective operations, and optimize data utilization through systematic experiments in data processing, model training, and evaluation with model performance as feedback. However, the continuous scaling of data and model sizes, coupled with the increasing complexity of processing operations, renders an exhaustive exploration of the combinatorial space of data recipes infeasible. Therefore, an essential question arises: can AI systems automatically generate a data recipe for training LLMs, including the orchestration of data pipelines and the implementation of each operation, in a cost-efficient way?
@@
To bridge this gap, we introduce a new task: end-to-end data recipe generation for LLM adaptation. As shown in Fig 1(a), given a target benchmark and a pool of available data sources, the objective is to generate a data recipe by specifying the precise data processing pipeline to yield training data for adapting an LLM to the target task. This task requires sophisticated reasoning to analyze heterogeneous data sources, apply domain-specific processing operations, and generate executable code. While reinforcement learning with verifiable rewards (RLVR) has proven effective in enhancing LLM reasoning in domains such as coding Zeng et al. (2025) and mathematics Yeo et al. (2025), applying this paradigm to our task poses two key challenges: (1) Data absence: As a novel task, there are no curated datasets or standardized evaluation benchmarks for data recipe generation. (2) Expensive and delayed supervision: While downstream performance naturally serves as the reward signal, it is impractical to incorporate full LLM training into an online RL loop.
@@
To address these challenges, we curate a comprehensive task pool comprising 31 widely used benchmarks across 19 distinct domains. These domains encompass reasoning-heavy fields, such as mathematics and coding, as well as knowledge-centric fields, such as finance, medicine, and natural sciences. The pool is partitioned into 25 training tasks and 6 held-out evaluation tasks, with each task supported by 8–15 source training datasets. We further propose a Data Verifier that estimates the utility of processed data without requiring model training, providing a low-latency proxy reward for scalable online RL. Empirical validation demonstrates that our Data Verifier correlates well with downstream performance and exhibits superior robustness across diverse tasks compared to existing data scoring metrics Li et al. (2024a); Friedman and Dieng (2023). Leveraging this task pool and proxy reward, we present DataChef-32B, an LLM specialized in generating optimal data recipes.
Extensive evaluations on 6 held-out tasks demonstrate that DataChef-32B matches the capabilities of the state-of-the-art proprietary model, Gemini-3-Pro, as illustrated in Fig. 1(b). Furthermore, recipes generated by DataChef-32B outperform the SOTA selection algorithm, DEITA Liu et al. (2024), on most tasks, highlighting the superior potential of exploring a vast coding space over relying on hand-designed selection heuristics. Notably, our recipes adapt Qwen3-1.7B-Base to achieve 66.7 on AIME’25 AIME (2025) and 46.3 on ClimaQA Manivannan et al. (2025), respectively, surpassing the official Qwen3-1.7B checkpoint with industry-level post-training on expert-curated data recipes.
@@ 2. $\bullet$
    We construct a large-scale and diverse data pool covering 19 domains, 31 benchmarks, and 257 datasets to facilitate research in this area.
@@ 3. $\bullet$
    We propose an efficient learning framework with a proxy reward that enables scalable online RL. Experiments show that DataChef-32B achieves performance comparable to top-tier proprietary models on the data recipe generation task.
@@ 3 Methodology
Figure 2: Illustration of task construction and DataChef framework. Given a task, a policy LLM generates a data recipe, which is executed to produce a training dataset. The Data Verifier then evaluates a sampled subset to provide a scalar reward, guiding the policy update via GRPO to optimize for data quality and executability.
@@ 3.2 Task Pool Construction
Seed Task Curation. We construct a diverse task pool encompassing 19 heterogeneous domains, including reasoning, coding, and knowledge-intensive fields such as healthcare, finance, and natural science. As shown in Fig 2(a), for each domain, we curate representative benchmarks (e.g., GSM8K and AIME’25 for mathematics), totaling 31 benchmarks. For each benchmark, we retrieve relevant candidate datasets from Hugging Face, prioritizing those with high community engagement (downloads and likes), yielding a repository of 257 distinct data sources. From this collection, we construct 25 seed tasks for training and reserve 6 held-out tasks (3 in-domain, 3 out-of-domain) for evaluation. Details of the selected benchmarks and automatic retrieval procedure for candidate datasets are provided in Appx. A.1.
@@ 3.3 End-to-end Data Recipe Generation
Cold-start Initialization. Training the policy from scratch using RL is non-trivial due to the low executability of data recipes, leading to sparse, high-variance rewards and ineffective exploration Shao et al. (2024); Liu et al. (2025c). To mitigate this, we employ a cold-start Supervised Fine-Tuning (SFT) phase. We observe that decoupling reasoning and coding yields superior inference-time performance (detailed in Sec. 4.4). Therefore, we construct a demonstration set using a decoupled generation process: a strong reasoning model proposes plans, and a specialized coding model implements them. Rejection sampling is used to retain only high-quality rollouts based on execution success and data quality. Initializing $\pi_{\phi}$ on this curated dataset equips the policy with a foundational capability for code generation, significantly stabilizing the subsequent RL phase.
Table 1: Main Results on six held-out tasks. We report the mean Data Verifier Score $\mathrm{DVS}_{\mathrm{avg}@32}$ and the Downstream Benchmark Score $\mathrm{DBS}$, where the Average column presents $\mathrm{DBS}_{\mathrm{norm}}$ as a normalized score relative to Single-Sourcebest (100.0). Qwen3-Next-80B $\oplus$ Kimi-K2 denotes a combination using Qwen3-Next-80B for reasoning and Kimi-K2-Instruct for coding. DataChef-32B achieves performance comparable to the closed-source Gemini-3-Pro and significantly outperforms other open-source baselines across all settings.
Note: All Methods are fine-tuned from Qwen3-1.7B-Base
@@
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}_{\mathrm{norm}}\uparrow$
Model Reference
Qwen3-1.7B-Base - 0.8 - 20.0 - 1.7 - 25.5
Qwen3-1.7B - 20.8 - 33.3 - 25.7 - 123.3
Human Preprocess $\oplus$ Data Selection
Single-Sourceavg - 6.1 - 23.4 - 6.3 - 63.9
Single-Sourcebest - 8.5 - 39.6 - 10.3 - 100.0
AlgorithmIFD - 7.0 - 3.3 - 4.6 - 45.3
AlgorithmDEITA - 7.5 - 6.7 - 10.9 - 70.5
LLM-as-a-Chef
@@
Qwen3-Next-80B $\oplus$ Kimi-K2 48.7 8.9 78.3 23.3 39.2 7.4 55.4 78.6
@@
$\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}\uparrow$ $\mathrm{DVS}_{\mathrm{avg}@32}\uparrow$ $\mathrm{DBS}_{\mathrm{norm}}\uparrow$
Model Reference
Qwen3-1.7B-Base - 15.1 - 33.7 - 14.2 - 35.9
Qwen3-1.7B - 44.2 - 73.4 - 59.8 - 101.0
Human Preprocess $\oplus$ Data Selection
Single-Sourceavg - 26.0 - 41.7 - 49.3 - 65.1
Single-Sourcebest - 43.6 - 63.7 - 70.3 - 100.0
AlgorithmIFD - 37.9 - 60.8 - 56.6 - 87.7
AlgorithmDEITA - 38.0 - 63.2 - 63.7 - 92.4
LLM-as-a-Chef
@@
Qwen3-Next-80B $\oplus$ Kimi-K2 41.5 42.6 54.7 64.0 8.6 4.1 34.9 68.0
@@ 5. $\bullet$
where $\lambda_{\emptyset}$ and $\lambda_{\mathrm{fmt}}$ are positive penalty coefficients. Please refer to Appx. A.2 for a detailed description of the category definitions used in the prompt and to Appx. E for computational cost analysis.
@@ 4.1 Setups
Figure 3: Correlation analysis of data evaluation metrics. (left) We summarize the Pearson correlation coefficients across all six evaluated tasks. (right) We detail the relationship between metric scores (X-axis) and downstream performance (Y-axis) on Language and Code tasks. The Data Verifier maintains a strong, consistent positive correlation across disparate domains. Please refer to Table 5 and Fig. 11 in Appx. D for complete results.
@@
Baselines. We compare DataChef-32B against leading LLMs, including Qwen3-32B, Kimi-K2-Instruct Team et al. (2025), Qwen3-Next-80B-A3B-Thinking, and Gemini-3-Pro Google (2025). Additionally, we incorporate the following results as reference: (1) To benchmark raw data quality and hand-designed strategies, we manually filter and format each available source, reporting the average and best downstream performance trained on single-source datasets. Furthermore, we apply SOTA data selection algorithms, IFD Li et al. (2024a) and DEITA Liu et al. (2024), to this manually pre-processed pool. (2) We report the performance of the official Qwen3-1.7B checkpoint, representing an industry-standard topline achieved via expert-curated data recipes. Detailed experiment setups are provided in Appx. B.
@@ 4.2 Main Results
Main Comparison. Table 1 presents the performance of DataChef-32B against baselines across in-domain and out-of-domain tasks. DataChef-32B achieves superior performance compared to a strong practical baseline, Qwen3-Next-80B $\oplus$ Kimi-K2, which leverages open-source state-of-the-art specialized models (Qwen3-Next-80B-A3B-Thinking for reasoning and Kimi-K2-Instruct for coding). Specifically, our end-to-end model surpasses this composite system with average improvements of +8.6% and +9.2% in $\mathrm{DVS}_{\mathrm{avg}@32}$, and +10.7% and +7.4% in $\mathrm{DBS}$ on in-domain and out-of-domain tasks, respectively. Notably, DataChef-32B achieves performance comparable to the closed-source top-tier Gemini-3-Pro, demonstrating exceptional robustness and effectiveness in automated data recipe generation.
@@
Surpassing Human Baselines. By selecting the most promising recipe from 32 samples (Oracle Upper Bound), DataChef-32B outperforms $\textsc{Single-Source}_{\mathrm{best}}$, IFD, and DEITA across most tasks, achieving an average in-domain score of 130.3. This indicates that DataChef goes beyond simple dataset selection and synthesizes novel data processing pipelines, including effective selection, mixing, synthesis, and filtering, thereby demonstrating the advantage of automatically exploring a vast code space over human-designed heuristics. Remarkably, it achieves 66.7 on AIME’25 and 46.3 on ClimaQA, surpassing the official checkpoint trained with industry-level, expert-curated data recipes. These results underscore the potential of fully automating data recipe generation for LLM training.